# Atomic Localization Fluorescent Microscopy

### v1

@@ Abstract
Super-resolution microscopy has revolutionized the imaging of complex physical and biological systems by surpassing the Abbe diffraction limit. Recent advancements, particular in single-molecular localization microscopy (SMLM), have pushed localization below nanometer precision [1, 2, 3, 4, 5], by applying prior knowledge of correlated fluorescence emission from single emitters [6]. However, achieving a refinement from $1$ nm to $1$ngström demands a hundred-fold increase in collected photon signal [7, 6, 8]. This quadratic resource scaling imposes a fundamental barrier in SMLM, where the intense photon collection is challenged by photo-bleaching, prolonged integration times, and inherent practical constraints [9]. Here, we break this limit by harnessing the periodic nature of the atomic lattice structure. Specifically, applying this discrete grid imaging technique (DIGIT) in a quantum emitter system, we observe an exponential collapse of localization uncertainty once surpassing the host crystal’s atomic lattice constant. We further applied DIGIT to a large-scale quantum emitter array, enabling parallel positioning of each emitter through wide-field imaging. These results showcase that DIGIT unlocks a potential avenue to applications ranging from identifying solid-state quantum memories in crystals to the direct observation of optical transitions in the electronic structure of molecules.
@@
Figure 1: DIGIT concept. a, Farfield PSF of an emitter with a diffraction-limited width $\sigma_{0}$ and its reconstructed location $(\mu_{x},\mu_{y})$ using SMLM. Multiple localizations’ distribution in one dimension in b, has a linewidth of $\sigma$. c, a zoomed-in view of the emitters’ underlying atomic structure - characterized by lattice constant a - illustrating their angström-level localization. d, transition of emitter localization from SMLM (purple) to DIGIT (green). e, DIGIT further extends in two dimensions. The color coding in SMLM represents distinct lattice groupings. f, The standard deviation of estimated location $\delta$: conventional SMLM in purple follows a polynomial scaling while DIGIT (green) scales exponentially as $\sigma$ falls below the lattice constant $a$
@@ I Introduction
Optical super-resolution microscopy resolves the position of fluorescent emitters with a precision $\sigma$, that scales with the number of collected photons $N$. Specifically, single-molecular localization microscopy (SMLM), which determines emitters’ positions by reconstructing their far-field point-spread functions (PSFs) [6], is fundamentally limited by $\sigma\geq\sigma_{0}/\sqrt{N}$, where $\sigma_{0}$ is the standard deviation of the PSF, as shown in Fig. 1a. Recent works have achieved sub-nanometer $\sigma$ localizing both fluorescent dyes [2, 4] and quantum emitters [10, 5, 11]. However, these approaches assume that independent measurements are made of a continuous position variable $\mu_{x}$ (Fig. 1b). In many physical systems, emitters reside at discrete lattice sites, prompting a breakdown of the continuous-space assumption. For example, crystal defects such as color centers are confined to specific atomic positions in the host lattice (Fig. 1c) [12, 13]. As localization approaches the lattice scale, it becomes essential to incorporate these positional constraints. In this work, we introduce a super-resolution technique that explicitly factors in the discrete nature of emitter positions, resulting in a new localization scaling of $\sigma\propto e^{-\sqrt{N}}$. This exponential enhancement pushes beyond the conventional limit, enabling localization approaching the atomic regime.
To illustrate this concept, Fig. 1d presents the transition from continuous to discrete localization. The purple dashed line represents the conventional SMLM localization distribution $f(\mu_{x})$. By quantitatively linking the emitters’ underlying atomic lattice structure (black lines), the posterior localization distribution $P(x)\propto f(\mu_{x})\sum\delta(x-na)$ becomes discretized, sharply peaking at each lattice site $n$ with a periodicity $a$ (green lines). A detailed explanation of this incorporation of lattice constraints within a Bayesian framework is provided in Supplementary Information Sec. I.
The approach can be generalized to two dimensions where the position of each emitter is expressed not as $\bm{\mu}_{i}\pm\sigma_{i}$ but as an integer combination of lattice vectors: $\bm{\mu}_{i}=m_{i}\mathbf{a}_{1}+n_{i}\mathbf{a}_{2}$, where $(\mathbf{a}_{1},\mathbf{a}_{2})$ are lattice unit vectors and $(m_{j},n_{j})\in\mathbb{Z}$. We call this implementation the discrete grid imaging technique (DIGIT), and Fig 1e illustrates the resulting discrete localizations. The key advantage of DIGIT, as shown in Fig 1f, is that after $\sigma$ approaches the lattice constant $a$, the standard deviation of emitters’ estimated localization $\delta$ diminishes exponentially, beyond the polynomial scaling in conventional SMLM.
In this paper, we first demonstrate lattice-scale localization precision and accuracy using a diamond sample by reengineering the SMLM system. We then extend it to achieve atomic-scale localization using the Bayesian-based DIGIT approach. Additionally, we apply DIGIT to $>600$ emitters and examine the enhancement requirements in a large-scale quantum system. Collectively, these advancements establish DIGIT as a competitive tool for attaining unprecedented precise measurements, ultimately paving the way to direct optical resolution of crystal and atomic features within quantum systems.
@@
Figure 2: Sub-diffraction imaging of SiVs. a, the schematic of sample preparation and characterization, including SiV FIB implantation and PL collection at cryogenic temperatures. b, Diffraction-limited PL image of implanted SiV clusters. c, PLE spectrum of the cluster highlighted in b, with Lorentzian fitting shown in green. The top inset displays an EMCCD image of resolved emitter 1 with scale bar of $200$ nm. d,e, Comparison of SMLM-reconstructed images without and with digital twin calibration. In d, closely spaced emitters appear unresolved, whereas in e, distinct emitters emerge with a localization uncertainty towards $\sigma=0.6$. f, The localization precision $\sigma$ improves following shot noise limited scaling.
II Lattice-scale Accuracy and Precision
Diamond color centers (CCs) serve as exemplary systems for DIGIT because of their long-lived emission and well-defined locations within the diamond crystal lattice. Although previous studies have achieved sub-nanometer localization $\sigma$ in diamond CCs [14, 15, 5, 1], several challenges remain in their imaging accuracy. These include mitigating the biases of the fitting model with fixed dipole nature of atomic transitions [16, 17, 18, 19] and drift-induced localization errors [20]. To address these challenges, we co-designed an end-to-end physical-digital twin system for SMLM, which accurately digitizes the experimental apparatus and optimizes its performance (Extended Data Fig. 2). By integrating this modeling into the microscope design and data acquisition, we achieve lattice-scale accuracy and precision.
The physical system involves a custom widefield cryogenic photoluminescence (PL) microscope (Fig. 2a) to image an array of silicon-vacancy centers (SiV) produced by focused-ion beam (FIB) implantation (See Methods). The farfield emission from SiVs is collected by an electron-multiplying charge-coupled device (EMCCD) camera. Fig. 2b shows the resulting PL image of SiV cluster arrays from off-resonant excitation.
To resolve individual SiV emitters beyond the diffraction limit, we applied prior knowledge of the natural inhomogeneous broadening of SiV transition energies [21] as illustrated in Fig. 2a. Specifically, we employed cryogenic photoluminescence excitation (PLE) by sweeping the excitation laser frequency across the zero-phonon line transitions and collecting light emitted into the phonon sideband on the EMCCD camera (see Methods). Fig. 2c presents the PLE spectrum of the cluster highlighted in Fig. 2b, revealing three distinct peaks with extracted resonance frequencies at $\omega_{1}/2\pi=406.708$ THz, $\omega_{2}/2\pi=406.724$ THz, $\omega_{3}/2\pi=406.735$ THz. With the laser frequency locked at each $\omega_{i}$, we subsequently re-acquired EMCCD images at extended integration times. For example, Fig. 2c shows an EMCCD image of SiV 1’s PSF with $\omega_{1}$.
Next, we developed a comprehensive digital twin model to extract CCs’ positions from their emission profiles. Instead of a simplified 2D Gaussian fit, our digital twin accurately represents the entirety of the physical apparatus, correcting for system-induced biases to achieve Ångström-level localization accuracy and precision. Built on a fixed-dipole model governing far-field dipole emissions [22, 23], it further incorporates imperfections of the PLE setup such as optical aberrations, interpixel non-uniformity, electronic noise in the EMCCD, and setup drift. We verified the model’s accuracy by examining two orthogonally polarized optical transitions within a single SiV [21] and observed the two distinct PSFs overlap. Further details are provided in Supplementary Information Sec. III.
Fig. 2d and e compare the resulting localization probability density function $f(\mu_{i})$ derived from (d) a basic 2D Gaussian fit and (e) our digital twin model. In fig. 2d, the simple Gaussian approach yields a localization uncertainty of $\sigma=1.8$ nm, insufficient to resolve two closely spaced emitters. However, by applying the accurate modeling, $\sigma$ reduces to $0.6$ with integrating PSF over $8000$ frames, as illustrated in Fig. 2e. Importantly, the three CCs are clearly spatially resolved. Fig. 2f plots the uncertainty $\sigma$ of the highlighted emitter in Fig. 2e as a function of integration time $t$. The standard error follows the shot-noise-limited scaling [24]. As $t$ increases, $\sigma$ decreases past the threshold at the diamond lattice constant $a$. The localization results for three emitters are provided in Supplementary Information Sec. IV.
@@ II Lattice-scale Accuracy and Precision
Figure 3: Atomic-scale precision achieved by DIGIT. a, Comparison of SMLM localization distribution $f(x)$ in purple versus DIGIT posterior distribution $P(x)$ in green, aligned along the lattice by grey dots. The inset illustrates the diamond lattice model projected along the [100] crystallographic direction. b, Standard deviation of the mean distance between all emitter pairs. The experimentally measured localization precision using DIGIT(green) with its simulation predictions(yellow), demonstrating an exponential improvement over the SMLM shot noise scaling.
III Atomic Localization with DIGIT
Figure 4: Large-scale DIGIT. a, Three-dimensional representation of widefield PLE obtained by sweeping a resonant laser over 30 GHz in a $10$ MHz increment. b, Reconstructed clusters, each color-coded by the number of emitters $M$. c, Histogram of measured precision $\sigma$ for individual resolved emitters whose $\sigma<0.3a$. d,e, DIGIT performance under varying cluster density conditions, demonstrated through numerical simulations (upper panels) and experimental results (lower panels). d studies DIGIT’s advantage over SMLM by evaluating $\delta_{P}/\delta_{f}$ at different emitter counts $M$ per cluster and e studies the effect of cluster size $L$ when $M=5$.
Building on the localization data from the previous section, we applied DIGIT to image SiVs with a Bayesian prior that encodes the diamond lattice coordinates. Since the PSF is collected along diamond’s [100] crystallographic direction, we simplified the intricate three-dimensional lattice by projecting it onto a two-dimensional plane along the [100] axis, as shown in Fig. 3a. We then computed $P(x)$ after fitting the lattice positions via an affine transformation whose parameters are optimized through maximum likelihood estimation (see Extended Data Fig. 3).
Fig. 3a illustrates the resulting lattice positions (grey dots) and the new SiV localization $P(x)$ determined by DIGIT. In contrast to the purple contour representing the conventional SMLM localization distribution $f(\mu)$, DIGIT localizations $P(x)$ (green dots) exhibit discrete peaks that align well with specific lattice sites, reflecting the Bayesian incorporation of diamond’s periodic lattice structure.
Fig. 3b quantifies DIGIT by looking at the average localization deviation $\delta(r)$ over three emitter pairs at different stages of $\sigma$. We compare the DIGIT results (green dots) with conventional SMLM represented (purple dashed line) and the Monte Carlo simulation of the DIGIT-based digital twin (yellow line). Our experimental result of $\delta(r)$ aligns closely with the DIGIT simulation, surpassing the conventional limit of SMLM. Error bars represent the confidence intervals of localization uncertainty. The inset of Fig. 3b zooms in on the regime where $\sigma<0.35a$ DIGIT enhancement becomes pronounced. Each green dot represents an individual measurement at a specific $\sigma$. As a result, DIGIT enables localization entering an atomic regime of $\delta_{P}(r)=0.178\pm 0.107$ below the diamond lattice spacing, while conventional SMLM yields a mean uncertainty of $\delta_{f}(r)=0.7$ for the same three emitter pairs. This 5-fold improvement over SMLM enables atomic resolution that would otherwise necessitate integration time 25 times longer.
IV Large-scale DIGIT
To explore the full capability of DIGIT, we extend the technique from one quantum cluster to multiple clusters across $20$x20202020 $\mu$m field of view. Using widefield PLE, we successfully resolved over $600$ single emitters within $200$ clusters. The localization of these emitters was facilitated by the pre-etched markers in diamond (Extended Data Fig. 1), which served as defined reference coordinates. Expanding to a wider field of view emphasizes the importance of the implementation of aberration corrections and camera interpixel non-uniformity calibrations to maintain high localization precision (see Supplementary Information Sec. III). Through the analysis of various clusters, we investigated the influence of parameters enabled by widefield measurements, particularly focusing on clusters’ density of states and brightness.

### v2

# A Bayesian Approach towards Atomically-precise Localization in Fluorescence Microscopy

@@ Abstract
Super-resolution microscopy has revolutionized the imaging of complex physical and biological systems by surpassing the Abbe diffraction limit. Recent advancements, particularly in single-molecule localization microscopy (SMLM), have pushed localization below nanometer precision, by applying prior knowledge of correlated fluorescence emission from single emitters. However, achieving a refinement from $1$ nm to $1$ Ångström demands a hundred-fold increase in collected photon signal. This quadratic resource scaling imposes a fundamental barrier in SMLM, where the intense photon collection is challenged by photo-bleaching, prolonged integration times, and inherent practical constraints. Here, we break this limit by harnessing the periodic nature of the atomic lattice structure. Applying this discrete grid imaging technique (DIGIT) in a quantum emitter system, we observe an exponential collapse of localization uncertainty once surpassing the host crystal’s atomic lattice constant. We further applied DIGIT to a large-scale quantum emitter array, enabling parallel positioning of each emitter through wide-field imaging. Collectively, these advancements establish DIGIT as a competitive tool for achieving unprecedented, precise measurements, ultimately paving the way to direct optical resolution of crystal and atomic features within quantum and biological systems.
@@
Figure 1: DIGIT concept. a, Farfield PSF of an emitter with a diffraction-limited width $\sigma_{0}$ and its reconstructed location $(\mu_{x},\mu_{y})$ using SMLM. b, Probability density function $f(\mu_{x})$ for a continuous variable $x$. c, A zoomed-in view reveals the emitters’ atomic structure, defined by a lattice with constant ($a_{1}$$a_{2}$). Unlike a continuous spatial distribution, the emitters’ positions are restricted to discrete sites determined by the underlying crystal lattice. d, Comparison of posterior localization probability $P(x)$ between conventional SMLM (purple) and DIGIT-equipped SMLM (green). e, An exemplary DIGIT localization in two dimensions. The color coding in SMLM represents distinct lattice groupings. f, Normalized localization precision $\sigma$ as a function of total detected number of photons NN: conventional SMLM follows the shot noise scaling (purple) while DIGIT exhibits exponential decay (green) when localization approaches lattice constant ($\sigma\sim a$).
@@ I Introduction
Optical super-resolution microscopy has become an important tool for fundamental studies in biology and physics due to its capability to resolve densely packed structures far below the diffraction limit. Recent advancements have enabled lateral resolutions into the sub-nanometer regime localizing fluorescent dyes [1, 2, 3, 4], nanoparticles [5] and quantum emitters [6, 7, 8]. However, pushing localization beyond sub-nanometer precision to atomic scales has remained challenging due to the fundamental shot noise scaling limit that scales inversely with the square root of the photon numbers $N$ [9, 10, 11]. This scaling implies that achieving finer spatial precision requires quadratically increased photon counts, quickly reaching practical limitations such as photon budgets, photobleaching, sample drift and slow noise processes [12]. These constraints significantly restrict real-time tracking of dynamic molecular processes.
The scaling limit arises due to the standard statistics of estimating the mean quantity (e.g. emitter position) over multiple independent measurements. This standard statistical procedure, however, does not capture the physical constraints that are present in many physical systems of interest. For instance, quantum defects reside at discrete lattice sites [13, 14], nuclear pores are organized by periodic protein rings [15, 16], and synthetic nanostructures follow designed grids [17, 18]. Here we introduce a discrete-grid imaging technique (DIGIT) that includes these physical constraints and transforms localization space from continuous to a discrete set of configurations. The localization process then becomes a decision between these emitter configurations, rather than independent estimation of each mean position. This enables the determination of emitter positions with resolution exponentially surpassing the conventional scaling.
We experimentally implement DIGIT to achieve atomic-scale super-resolution imaging of quantum emitters in diamond. Quantum emitters have been extensively studied over the past decade, emerging as a leading platform for quantum sensing [19, 20, 21] and quantum networking [22, 23]. Previous experiments have achieved resolution around $\sigma\sim 0.7$ nm due to scaling limits inherent in conventional localization methods [24, 25, 26]. In this work, we apply widefield photoluminescence excitation (PLE) with DIGIT to overcome the shot-noise scaling of localization with photon counts, achieving an unprecedented localization precision of 0.178 Å. We further demonstrate that this method can be extended to large-scale emitter arrays, enabling massively parallel localization for different emitter clusters. These results show that DIGIT unlocks a potential avenue to applications ranging from identifying solid-state quantum memories in crystals to the direct observation of optical transitions in the electronic structure of molecules.
II Results
II.1 DIGIT concept
We introduce the DIGIT in the context of single-molecule localization microscopy (SMLM). SMLM is one of the most powerful super-resolution modalities, enabling the visualization of emitters down to the molecular level [1]. In conventional SMLM, each emitter produces a diffraction-limited point-spread function PSF, whose centroid $(\mu_{x},\mu_{y})$ is estimated with precision $\sigma$, fundamentally limited by $\sigma\geq\sigma_{0}/\sqrt{N}$, where $\sigma_{0}$ is the standard deviation of the PSF and $N$ is the total photon count [9] (Fig. 1a). After multiple SMLM measurements, we obtain a continuous probability density function of emitter locations $f(\mu_{x})$, as shown in Fig. 1b. In the case of solid-state crystal, however, emitters are only allowed to sit at discrete lattice sites identified by the substitute carbon atoms (Fig. 1c). In our implementation, we account for this by using a Bayesian statistical process, where we update the initial probability distribution $f(\mu_{x})$ to the posterior localization $P(x)$ given this underlying atomic lattice structure $\delta(x-na)$, where $a$ is the lattice constant. This posterior collapses onto sharp peaks at $x=na$, as depicted in Fig. 1d. This concept naturally extends to two dimensions as shown in Fig. 1e, where each emitter position is mapped onto the lattice grid as a discrete integer combination of lattice vectors: $\bm{\mu}_{i}=m_{i}\mathbf{a}_{1}+n_{i}\mathbf{a}_{2}$, with $(\mathbf{a}_{1},\mathbf{a}_{2})$ denoting lattice unit vectors and $(m_{j},n_{j})\in\mathbb{Z}$. The color-coded SMLM localization illustrates how DIGIT refines each localization onto the lattice. The resulting DIGIT localization $\sigma_{p}$ is shown in Fig. 1f as a function of $N$. Compared with the conventional position estimation, $\sigma_{p}$ represented in the green line scales $\sigma_{p}\propto e^{-\sqrt{N}}$, diminishing exponentially after $\sigma$ surpasses the lattice constant $a$, achieving atomic localization.
@@ II.1 DIGIT concept
Figure 2: Sub-diffraction imaging of SiVs. a, Schematic of sample preparation and characterization. SiVs are created via FIB implantation. The quantum emitters are optically excited using a 532 nm green laser, and the resulting PL is collected through a confocal microscope operated at cryogenic temperatures. Inset: Energy level structure of SiV centers. The optical transition frequency $\omega$ exhibits inhomogeneous broadening across different emitters, enabling SMLM with PLE spectroscopy. b, Diffraction-limited PL image of implanted SiV clusters. c, Top: EMCCD image of one resolved emitter. Scale bar: $200$ nm. Down: PLE spectrum of the cluster. Three emitters separate in frequency space within the diffraction-limited spot, with Lorentzian fitting shown in green. d, e, Comparison of SMLM-reconstructed images without and with digital twin calibration. In d, closely spaced emitters appear unresolved, whereas in e, distinct emitters emerge with a localization uncertainty towards $\sigma=0.6$ Å. f, The localization precision $\sigma$ improves following the photon-shot-noise-limited scaling with an SMLM technique.
II.2 SMLM for a quantum emitter
A prerequisite for implementing DIGIT is achieving localization accuracy $\sigma$ comparable to the lattice constant $a$. Here, we adapted PLE to achieve lattice-scale accuracy and precision. Compared with other emitter localization methods like stochastic optical reconstruction microscopy [6, 27], structured illumination microscopy [28, 8, 29] and quantum state control [30, 31], PLE allows deterministic excitation control of individual targeted emitters in a dense cluster by spectrum separation due to the inhomogeneous distribution nature of quantum emitters.
We use silicon-vacancy (SiV) centers as exemplary quantum emitters due to their exceptional stability and resistance to photobleaching, making them well-suited for repeated or prolonged imaging applications. Moreover, the mature focused ion beam (FIB) enables nanometer-scale resolution in the implantation process. Despite the high spatial precision, the overall number of implanted centers can be kept low, which helps to maintain well-isolated emission sites and minimizes background interference.
We first use PLE spectroscopy to spectrally resolve individual emitters. As shown in Fig. 2a, a SiV array is implanted in diamond with FIB and imaged with a widefield cryogenic microscope. A representative photoluminescence (PL) image is displayed in Fig. 2b, where each bright spot may contain multiple SiVs separated by sub-diffraction-limit distances. To resolve these closely spaced emitters individually, we sweep the excitation laser frequency across the zero-phonon-line transitions while simultaneously recording emission into the phonon sideband using an electron-multiplying charge-coupled device (EMCCD) camera. As shown in Fig. 2c, three distinct spectral peaks at frequencies $\omega_{1},~\omega_{2},~\omega_{3}$ correspond to three individual emitters within the cluster shown in Fig. 2b.
@@ II.2 SMLM for a quantum emitter
Figure 3: Atomic-scale precision achieved by DIGIT. a, Comparison of SMLM localization distribution $f(x)$ in purple versus DIGIT posterior distribution $P(x)$ in green, aligned along the lattice by grey dots. The inset illustrates the diamond lattice model projected along the [100] crystallographic direction, including lattice’s global rotation $\theta$ and offset $U$. b, Normalized localization precision $\sigma$ distance between all emitter pairs as a function of total collected number of photons NN. The experimentally measured localization precision using DIGIT (green) with its simulation predictions (yellow), demonstrating an exponential improvement over the SMLM shot noise scaling (purple). The inset zooms in the transition where DIGIT localization begins to surpass conventional SMLM.
We then sequentially localize these three emitters with multiple localizations. The initial localization uncertainty fitted by a 2D Gaussian distribution yields $\sigma=1.85$ nm over $8000$ frames, as shown in Fig. 2d. This precision is insufficient to resolve the individual emitters due to the inherent inaccuracies in the Gaussian approximation for the collected PSFs. To address this limitation, we further apply established calibration and correction techniques tailored to the emitters’ system. These include EMCCD inter-pixel gain correction, optical aberration correction through adaptive optics and PSF modeling, dipole-orientation characterization, and drift correction via Kalman filter [3, 32]. These improvements enable us to achieve a localization precision $\sigma=0.6$ Å as illustrated in Fig. 2e. The localization precision follows the shot-noise-limited scaling (Fig. 2f) [33] and surpasses the threshold at the diamond lattice constant $a$ when t>300t>300~s.
II.3 Atomic Localization with DIGIT
Figure 4: Large-scale DIGIT. a, Three-dimensional representation of widefield PLE obtained by sweeping a resonant laser over 30 GHz in a $10$ MHz increment. b, Number of emitters in each reconstructed cluster. c, Histogram of the measured precision $\sigma$ for individually resolved emitters. d, Top: Numerical simulations of the precision ratio $\sigma_{p}/\sigma_{f}$ as a function of the number of measured photons NN for different numbers of emitters $M$. Bottom: Representative experimental data with theoretical fits for $M=3,4,5$. Dashed line: $\sigma_{p}/\sigma_{f}=1$. e, Top: Numerical simulations of the precision ratio $\sigma_{p}/\sigma_{f}$ as a function of the number of measured photons NN for various cluster sizes $L$, with $M=5$ fixed. Bottom: Representative experimental data with theoretical fits showing the dependence on $L\geq 3000$ and $L\leq 50$. Dashed line: $\sigma_{p}/\sigma_{f}=1$.
Building upon the localization results, we applied DIGIT to map the positions of SiVs, encoding their locations with respect to the diamond lattice constraints. Although the full three-dimensional diamond lattice is inherently complex, we simplify it by projecting the lattice along the [100] direction of the diamond substrate $\delta(x,y)$ [34], as shown in Fig. 3a. To further enhance the accuracy of the Bayesian model used in DIGIT, we mapped the measured coordinates to lattice sites through an affine transformation. This transformation accounts for the global rotation $\theta$ and offset of the lattice $U$, aligning the experimental data with the periodic structure of the lattice. The parameters for this transformation were optimized using maximum-likelihood estimation, as detailed in Extended Data Fig. 3. The resulting DIGIT localization of the three emitter clusters is shown in Fig. 3a, where the conventional SMLM localization distribution breaks into discrete peaks that align with the fitted lattice sites, significantly enhancing the localization probability at certain lattice sites.
We then analyze the localization scaling as a function of $N$. With a conventional SMLM, the localization is proportional to the square root of the measured photon number $N$, as shown of the purple line in Fig. 3b. By applying the concept of DIGIT, when the precision $\sigma$ is much larger than the lattice constant ($\sigma\gg a$), the precision follows the same trajectory as the conventional SMLM. However, when $\sigma/a$ approaches $0.33$, the localization scaling deviates from the conventional scaling and drops exponentially with the number of measured photons, as depicted in the yellow curve. We then demonstrate this concept with experimental data. As shown in Fig. 3b, the experimental results of the emitter localization fit well with the modeling and demonstrate an unprecedented localization of $\sigma_{p}=0.178\pm 0.107$ Å below the diamond lattice spacing, while conventional SMLM yields a mean uncertainty of $\sigma_{f}=0.7$ Å for the same three emitter pairs.
II.4 Large-scale DIGIT
[…]